# Aligning Language Models with Real-time Knowledge Editing

### v2

@@ 4.4 Self-adaptive Post-alignment Inference
In the inference phase, we employ both the original pre-alignment model and the post-alignment adapter, which does not consume extra memory while enabling two inference paths. Each time a query $q$ is requested, the filter-based smart retriever is utilized to recall possible relevant edits in the memory. If there exists a set of relevant edits $\mathcal{E}^{*}$, indicating that the query involves edited knowledge, we fill both $\mathcal{E}^{*}$ and $q$ into the knowledge editing prompt (see Section 2.2) and then feed the prompt into the model with the post-alignment adapters activated. Otherwise, the query directly goes through the pre-alignment model without any processing. The inference strategy is formularized below:
@@
where $a$ denotes the final answer, $f_{\Phi}$ denotes the original model parametrized by $\Phi$, $f_{\Phi+\Delta\Phi(\Theta)}$ denotes the model with post-alignment adapters, and $\mathbf{KEPrompt}(\cdot,\cdot)$ denotes the knowledge editing prompt template.
Method CRAFT-Statistics CRAFT-Finance
@@
KEDAS (ours) 99.80 100.00 100.00 81.20 62.17 99.85 62.50 100.00 55.95 70.61
@@ 5 Experiments
We evaluate representative knowledge editing methods and our proposed KEDAS on CRAFT. Our experiments are based on the framework of EasyEdit Wang et al. (2024b).
@@ 5.1 Experimental Setup
Language Model
In our experiments, we focus on Llama-3.1-8B-Instruct Llama Team (2024), a widely used open-source language model.
Evaluated Methods
We evaluate the following knowledge editing methods:
  • LoRA Zhang et al. (2024) directly updates knowledge by LoRA-based finetuning.
  • ROME Meng et al. (2022) locates multilayer perceptron (MLP) weights of related knowledge and writes in new key-value pairs.
  • IKE Zheng et al. (2023) leverages in-context examples of copying, updating and retaining knowledge to edit knowledge via prompting. Due to limited resources, we only evaluate the 8-shot setting.
  • EREN Chen et al. (2024b) prompts LLMs to decide the relevance of retrieved edits and then queries LLMs with or without edits.
  • WISE Wang et al. (2024a) stores edits in a parametric side memory module and utilized a router to decide which memory to go through.
  • LTE Jiang et al. (2024) aligns LLMs with the knowledge editing task by post-training. We adopt paraphrase-multilingual-MiniLM-L12-v2 Reimers and Gurevych (2019) as the retriever and set the number of adopted edits $k$ as 3, the default setting of LTE.
  • KEDAS is our proposed method. We use the same retriever as LTE and adopt bert-base-chinese Devlin et al. (2019) as the binary filter. We set the number of retrieved candidates $n$ as 8 and that of adopted edits $k$ as 3, in line with LTE for fair comparison.
Evaluation Metrics
@@
5.2 Main Results
The main results on CRAFT are presented in Table 5. Previous methods of knowledge editing show limitations in certain aspects. LoRA and ROME, as traditional methods primarily designed for single editing or limited number of edits, show no advantage when applied to the sequential editing setting on CRAFT. IKE, the retrieval-based approach, suffers from poor locality and mediocre portability, albeit achieves the highest edit success scores. The prompting-based EREN requires specific prompt optimization for different LLMs and can scarcely perform editing with its official default prompt, as can be seen from the low edit success. WISE is parameter-based and suffers from downgrading performance under the sequential setting, with weak performance in most metrics. LTE, with an alignment stage, achieves relatively robust performance in edit success and portability but struggles to achieve locality due to its fixed inference path. These results indicate that CRAFT is a challenging benchmark for existing knowledge editing methods, making it hard for them to achieve a consistently balanced performance.
@@ 5.2 Main Results
In Appendix G, we also compare KEDAS with previous methods on traditional knowledge editing benchmarks. The results show that KEDAS significantly outperform previous methods with balanced performance and confirm the effectiveness of KEDAS even on traditional benchmarks.
@@
Table 6: Ablation results. The highest scores are bolded.
5.3 Ablation Study
@@ 6 Related Work
6.1 Knowledge Editing Benchmarks
Conventionally, knowledge editing benchmarks are created based on temporal changes or counter-fact modifications of world knowledge either from existing QA datasets Levy et al. (2017); Yao et al. (2023) or Wikipedia Cohen et al. (2024a); Zhong et al. (2023), in which Zhong et al. (2023) first propose multi-hop composite portability that involves multiple edits. However, these are all static and suffer from data contamination as time goes by.
Thede et al. (2025) propose WikiBigEdit, a dynamic benchmark based on periodical changes in Wikipedia. However, only a small fraction (3.6%) of its samples support reasoning portability evaluation, limiting its comprehensiveness. Similarly, Tang et al. (2025) propose EvoWiki, a Wiki-based real-time dataset (primarily designed for the RAG), where each instance is associated with reasoning portability tests. Although EvoWiki claims to automatically update with new edits, its implementation and data pipeline remain unpublished.
Besides conventional knowledge editing benchmarks, there have also been some benchmarks designed for special use. Mitchell et al. (2022b) propose ConvSent to edit LLMs’ sentiments on given topics. Ishibashi and Shimodaira (2024) propose Sanitation to forget specific information stored in LLMs and thus address privacy concerns. Huang et al. (2025) propose HalluEditBench to evaluate whether knowledge editing correct hallucinations.
Our proposed CRAFT benchmark leverages statistical and financial data, which are both real-time and inherently dynamic. The strong interdependence among indicators allows each pair of edits to correspond to a meaningful reasoning portability test. This enables comprehensive evaluation across five dimensions: benchmark type, real-time property, real-world grounding, reasoning portability, and composite reasoning portability.
@@ 6.2 Knowledge Editing Methods
This line of work aims to edit model parameters or add extra parameters to perform knowledge editing. Meng et al. (2022)Meng et al. (2023) and Fang et al. (2025) adopt a locate-then-edit manner and aim to precisely edit MLP weights. Mitchell et al. (2022a) adopts meta-learning Hospedales et al. (2022) to predict the change of model parameters. Dong et al. (2022)Huang et al. (2023) and Wang et al. (2024a) add extra parameters to store edits in a parametric form. Mitchell et al. (2022b) trains a separate counter-fact model and does not involve retrieval (thus categorized as parameter-based), failing to fully exploit the capability of LLMs themselves.
@@
This line of work stores edits in a memory module and retrieve from it when needed. Zheng et al. (2023) leverages in-context examples of copying, updating and retaining knowledge to edit knowledge via prompting. Chen et al. (2024b) first prompts LLMs to decide the relevance of retrieved edits and then queries LLMs with or without in-context edits. Chen et al. (2024a) converts edits into soft prompt tokens with trained models and then concatenates user prompts to modify the behavior of frozen LLMs. Jiang et al. (2024) aligns LLMs with knowledge editing by post-training with in-context edits. Our proposed KEDAS can also be categorized as retrieval-based.
@@ 7 Conclusion
In this work, we introduce CRAFT, a novel automatically updatable real-time benchmark for knowledge editing that focuses on national statistics and finance of China. We evaluate representative knowledge editing methods on CRAFT and reveal that existing methods have certain limitations. We then propose a simple yet effective approach, KEDAS, to better align LLMs with real-time knowledge editing. Experiments validate the effectiveness of KEDAS on CRAFT in all aspects including edit success, locality and portability.
@@
First, the CRAFT benchmark focuses on the Chinese language only (but our methodology is universal and can be applied to any other languages) and is limited to the domains of national statistics and finance (but it is easily expandable to domains that data are periodically updated and automatically accessible). Second, the composite reasoning portability of CRAFT is merely 2-hop (but can be expanded to multiple hops via combination of time and indicators). Third, the Finance subset of CRAFT can be updated annually only (but can be improved to be quarterly updated by substituting the annual financial statements to quarter ones). Fourth, only one Llama-3 LLM is employed due to limited time and a limited set of existing knowledge editing methods is evaluated in our experiment due to limited resources and compatibility issues.
@@
All the experiments are fully reproducible since all the methods are deterministic and sampling is disabled during LLM generation.
@@
paraphrase-multilingual-MiniLM-L12-v2 Apache
EasyEdit MIT
Llama-3 Llama 3 Community License Agreement
@@ Appendix A Details of Alignment Data for KEDAS
In this way, the LLM learns to leverage retrieved edits for relevant queries while keeping the answer unchanged for locality queries.
Appendix B Details of LLM Alignment
@@
Table 8: Hyper-parameters for LLM finetuning.
We adopt LLaMA-Factory Zheng et al. (2024) to finetune the LLM. The alignment is done on an NVIDIA A40 48GB GPU and takes 2 hours.
The main training hyperparameters are presented in Table 8, most remain unchanged as the default values. Note that we train the LLM for only one epoch because finetuning for multiple epoches will cause over-fitting.
@@ Appendix C Details of Diverse Edit Augmentation
Figure 2: Prompt template for the declarative form.
Figure 3: Prompt template for the aliased form.
@@ Appendix D Details of Training Data for the Filter
Appendix E Example from CRAFT-Statistics
@@
Appendix F Example from CRAFT-Finance
@@
Appendix G Results on Traditional Knowledge Editing Benchmarks

### v3

@@ 4.4 Self-adaptive Post-alignment Inference
In the inference phase, we employ both the original pre-alignment model and the post-alignment adapter, which does not consume extra memory while enabling two inference paths. Each time a query $q$ is requested, the filter-based smart retriever is utilized to recall possible relevant edits in the memory. If there exists a set of relevant edits $\mathcal{E}^{*}$, indicating that the query may involve edited knowledge, we fill both $\mathcal{E}^{*}$ and $q$ into the knowledge editing prompt (see § 2.3) and then pass the prompt to the model with the post-alignment adapters activated. Otherwise, the query directly goes through the pre-alignment model without any processing. The inference strategy is formularized below:
@@
where $a$ denotes the final answer, $f_{\Phi}$ denotes the original model parametrized by $\Phi$, $f_{\Phi+\Delta\Phi(\Theta)}$ denotes the model with post-alignment adapters $\Delta\Phi(\Theta)$, and $\mathbf{KEPrompt}(\cdot,\cdot)$ denotes the knowledge editing prompt template.
4.5 Discussion
KEDAS borrows the idea of inference-time retrieval from EREN Chen et al. (2024b) and the idea of knowledge editing alignment from LTE Jiang et al. (2024). The novelty of KEDAS lies in the two unique designs: diverse edit augmentation and filter-based smart retriever paired with self-adaptive post-alignment inference. The former follows the idea that edits can go beyond their initial forms, while the latter demonstrates the insight that we do not have to perform knowledge editing for irrelevant queries and a lightweight flexible mechanism leads to more balanced performance.
Model Method CRAFT-Statistics CRAFT-Finance
@@ 4.5 Discussion
KEDAS 99.80 100.00 100.00 81.20 62.17 99.85 62.50 100.00 55.95 70.61
Qwen2.5 LoRA 43.81 18.86 0.00 39.48 35.91 24.73 11.80 0.00 31.59 32.89
ROME 35.06 32.79 0.00 29.00 26.90 9.75 2.95 0.00 14.39 9.34
IKE 96.82 19.44 82.40 48.48 42.02 97.96 28.78 81.80 39.27 46.16
EREN 28.32 61.08 92.60 34.55 30.67 31.25 42.43 93.00 33.69 29.53
WISE 40.72 43.46 94.40 34.57 38.39 28.07 51.43 95.00 28.67 21.25
LTE 99.95 8.38 6.60 77.67 58.92 99.96 31.56 5.20 51.86 60.18
KEDAS 99.95 100.00 100.00 80.88 58.59 99.88 66.47 100.00 51.55 65.97
@@ 5.1 Experimental Setup
5.1.1 Evaluation
@@
In the main experiments, we focus on sequential editing for both evaluation datasets.
5.1.2 Language Models
For CRAFT, we use Llama-3.1-8B-Instruct Llama Team (2024) and Qwen2.5-3B-Instruct Qwen (2025), both from widely used herds of open-source LMs (Llama-3.1 and Qwen2.5) and with different scales. For KnowEdit, we focus on the same Llama-3.1 model and provide the results on two other LMs in Appendix I.
5.1.3 Evaluated Methods
We evaluate the following knowledge editing methods: (1) LoRA Zhang et al. (2024) directly updates knowledge by LoRA-based finetuning. (2) ROME Meng et al. (2022) locates multilayer perceptron weights of related knowledge and writes in new key-value pairs. (3) IKE Zheng et al. (2023) leverages in-context examples of copying, updating and retaining knowledge to edit knowledge via prompting. (4) EREN Chen et al. (2024b) prompts LMs to decide the relevance of retrieved edits and then queries LMs with or without edits. (5) WISE Wang et al. (2024a) stores edits in a parametric side memory module and utilized a router to decide which memory to go through. (6) LTE Jiang et al. (2024) aligns LMs with the knowledge editing task by post-training. (7) KEDAS is our proposed method.
Due to compatibility, cost, and performance issues, we do not include some other baselines as discussed in Appendix F.
5.1.4 Implementation Details
Our experiments are based on the framework of EasyEdit Wang et al. (2024b). For most of the baselines, we follow the default hyper-parameter settings provided by EasyEdit.
For IKE, we only evaluate the 8-shot setting because of limited resources. For both LTE and KEDAS, we use paraphrase-multilingual-MiniLM-L12-v2 Reimers and Gurevych (2019) as the retriever and set the number of adopted edits $k$ as 3, the default setting of LTE. For KEDAS, we adopt bert-base-chinese Devlin et al. (2019) as the binary filter, which is trained based on the training set of CRAFT with details specified in Appendix D. We set the number of retrieved candidates $n$ as 8 and that of adopted edits $k$ as 3, in line with LTE. The diverse edit augmentation of KEDAS is done by prompting gpt-4o-mini [^7] https://platform.openai.com and the prompts are in Appendix C.
5.2 Results on CRAFT
The results on CRAFT are presented in Table 5.
Previous methods of knowledge editing show limitations in certain aspects. LoRA and ROME, as traditional methods primarily designed for single editing or a limited number of edits, show no advantage when applied to the sequential editing setting on CRAFT. IKE, the retrieval-based approach, suffers from poor locality and mediocre portability, albeit achieves outstanding edit success. The prompting-based EREN requires specific prompt optimization for different LMs and can scarcely perform editing with its default prompt, as can be seen from the low edit success. WISE is parameter-based and suffers from downgrading performance under the sequential setting, with weak performance in most metrics. LTE, with an alignment stage, achieves relatively robust performance in edit success and portability but struggles to achieve locality due to its fixed inference path. These indicate that CRAFT is challenging for existing knowledge editing methods, making it hard to achieve a consistently balanced performance.
[…]
5.3 Results on KnowEdit
Method ZsRE WikiBio $_{\text{recent}} WikiData$$_{\text{counteract}}$
ES L P HM ES L HM ES L P HM ES L P HM
Table 6: Results on KnowEdit. Metrics include edit success (ES), locality (L), portability (P) and harmonic mean (HM) of the previous metrics. The highest scores are bolded.
The results on KnowEdit are shown in Table 6. KEDAS exhibits outstanding performance in all the metrics and secures the highest harmonic mean (HM) scores across all the datasets. On average, its HM outperforms WISE by +26.9 and surpasses LTE by +18.2. As an alignment-based approach, KEDAS is superior especially in locality compared with LTE, with an improvement of +31.7.
On KnowEdit, we also further experiment with other LMs and under the traditional setting of single editing. These results can be found in Appendix I, where KEDAS also secures consistent advantage with balanced performance.
5.4 Ablation Study
@@
Table 7: Ablation results on CRAFT. The highest scores are bolded.
@@ 6 Related Work
6.1 Related Evaluation Datasets
Conventionally, knowledge editing datasets are created based on temporal changes or counter-fact modifications of world knowledge either from existing QA datasets Levy et al. (2017); Yao et al. (2023) or Wikipedia Cohen et al. (2024a); Zhong et al. (2023), in which Zhong et al. (2023) first propose multi-hop composite portability. However, these are all static and suffer from data contamination as time goes by.
Thede et al. (2025) propose WikiBigEdit, a dynamic dataset based on periodical changes in Wikipedia. However, only a small fraction (3.6%) of its samples support reasoning portability evaluation, limiting its comprehensiveness. Similarly, Tang et al. (2025) propose EvoWiki, a Wiki-based real-time dataset primarily designed for RAG. Although EvoWiki claims to automatically update with new edits, its implementation and data pipeline remain unpublished.
Besides conventional knowledge editing datasets, there have also been some designed for special use. Mitchell et al. (2022b) propose ConvSent to edit LMs’ sentiments on given topics. Ishibashi and Shimodaira (2024) propose Sanitation to forget specific information stored in LMs and thus address privacy concerns. Huang et al. (2025) propose HalluEditBench to evaluate whether knowledge editing correct hallucinations.
Our proposed CRAFT leverages statistical and financial data, which are both real-time and inherently dynamic. The strong interdependence among indicators allows each pair of edits to correspond to a meaningful reasoning portability test. This enables comprehensive evaluation across five dimensions: data type, real-time property, real-world grounding, reasoning portability, and composite reasoning portability.
@@ 6.2 Knowledge Editing Methods
This line of work edits model parameters or adds extra parameters for knowledge editing. Meng et al. (2022), Meng et al. (2023) and Fang et al. (2025) adopt a locate-then-edit manner and aim to precisely edit MLP weights. Mitchell et al. (2022a) adopt meta-learning Hospedales et al. (2022) to predict the change of model parameters. Dong et al. (2022), Huang et al. (2023) and Wang et al. (2024a) add extra parameters to store edits in a parametric form. Mitchell et al. (2022b) train a separate counter-fact model and does not involve retrieval.
@@
This line of work stores edits in a memory module and retrieve from it when needed. Zheng et al. (2023) leverage in-context examples of copying, updating and retaining knowledge to edit knowledge via prompting. Chen et al. (2024b) first prompt LMs to decide the relevance of retrieved edits and then query LMs with or without in-context edits, which may seem conceptually related to KEDAS but is essentially different in that it relies merely on prompting. Chen et al. (2024a) convert edits into soft prompt tokens with trained models and then concatenate user prompts to modify the behavior of frozen LMs. Jiang et al. (2024) align LMs with knowledge editing by post-training with in-context edits. Our proposed KEDAS can also be categorized as retrieval-based.
@@ 7 Conclusion
We have introduced CRAFT, a novel automatically updatable real-time dataset for knowledge editing that focuses on national statistics and finance of China. We evaluate representative knowledge editing methods on CRAFT and reveal that existing methods have certain limitations. We then propose a simple yet effective approach, KEDAS, to better align LMs with real-time knowledge editing. Experiments validate the effectiveness of KEDAS on CRAFT in all aspects including edit success, locality and portability. We hope this work may serve as a catalyst for shifting the focus of knowledge editing from static update to dynamic evolution.
Acknowledgments
This work is supported by the National Natural Science Foundation of China (62076008).
@@ Acknowledgments
Domain
CRAFT focuses on the Chinese language only and is limited to the domains of national statistics and finance. Nevertheless, our methodology is universal and can be applied to any other languages and CRAFT is easily expandable to domains that data are periodically updated and automatically accessible.
Portability
The composite reasoning portability of CRAFT is merely 2-hop. Nevertheless, it can be expanded to multiple hops via combination of time and indicators.
Baselines
Only a limited set of existing methods is evaluated in our experiments due to limited resources and compatibility issues.
Filter
The performance of KEDAS heavily depends on the filter. If the filter is not sufficiently robust, the overall performance of KEDAS will also be less competitive.
@@
Our work is fully reproducible since we have provided our implementation details and scripts for experiments.
@@
MiniLM Apache
EasyEdit & KnowEdit MIT
Llama-3 Llama 3 Community License
Qwen2.5 Apache
Llama-2 Llama 2 Community License
@@ Appendix A Details of Alignment Data for KEDAS
In this way, the LM learns to leverage retrieved edits for relevant queries while keeping the answer unchanged for locality queries.
Appendix B Details of LM Alignment
@@
Table 9: Hyper-parameters for LM finetuning.
We adopt LLaMA-Factory Zheng et al. (2024) to finetune the LM. The alignment is done on an NVIDIA A40 48GB GPU and takes 2 hours.
The main training hyperparameters are presented in Table 9, most remain unchanged as the default values. Note that we train the LM for only one epoch because finetuning for multiple epoches will cause over-fitting.
Prompt Template for Declarative You are a helpful assistant. You are given a query and a target new. Please generate the declaration form of the query and target new. Here is an example: Query: 2024年华大智造的总资产(亿元)是多少? Target new: 103.15 Declaration: 2024年华大智造的总资产(亿元)是103.15。 Please generate the declaration form of the query and target new below: Query: {query} Target new: {target_new} Declaration:
Figure 2: Prompt template for the declarative form.
@@ Appendix D Details of Training Data for the Filter
Appendix E Details Regarding KnowEdit
Dataset ZsRE WikiBio $_{\text{recent}} WikiData$$_{\text{counterfact}}$
# Train 10000 464 570 1428
# Test 1304 306 1266 885
Table 10: Data statistics of KnowEdit.
Table 10 presents the data statistics of KnowEdit.
Following Zhang et al. (2024) and Jiang et al. (2024), our evaluation metrics for KnowEdit include edit success (ES), locality (L) and portability (P). We also calculate the harmonic mean (HM) of the three scores as the edit quality. Note that WikiBio does not contain portability queries and we thus calculate the harmonic mean based on the first two metrics.
Appendix F Other Baselines
In our primary experiments, we also tried to evaluate many other modern baselines but encountered different sorts of issues (including compatibility, availability, cost, performance, etc.).
[…]
MEND Mitchell et al. (2022a), SERAC Mitchell et al. (2022b) and RECIPE Chen et al. (2024a) require a lot of resources that we cannot afford currently for training and do not provide trained checkpoints for the Llama-3 herd of LMs used in our experiments. Therefore, we do not include these in our main experiments.
Nevertheless, for a comprehensive comparison, we experiment on KnowEdit with all the mentioned baselines under the single editing setting using Llama-2-7B-Chat Touvron et al. (2023) due to that all the methods are compatible with single editing and the Llama-2 herd of LMs. The results are presented in Appendix I.
Prompt Template for Aliased You are a helpful assistant. You are given a query, a target new, and a declaration. Please generate a paraphrased sentence of the declaration with the central term translated to English. Here is an example: Query: 2024年华大智造的总资产(亿元)是多少? Target new: 103.15 Declaration: 2024年华大智造的总资产(亿元)是103.15。 Paraphrased sentence: 2024年华大智造的Total Assets(100 million yuan)是103.15。 Here is another example: Query: 2024年9月中国的订销报纸份数当期值(万份)是多少? Target new: 137885.2 Declaration: 2024年9月中国的订销报纸份数当期值(万份)是137885.2。 Paraphrased sentence: 2024年9月中国的Issue of Newspapers, Current Period Value(10000 pieces)是137885.2。 Please generate a paraphrased sentence of the declaration below: Query: {query} Target new: {target_new} Declaration: {declaration} Paraphrased sentence:
Figure 3: Prompt template for the aliased form.
Appendix G Example from CRAFT-Statistics
@@
Appendix H Example from CRAFT-Finance
@@
Appendix I More Results on KnowEdit
Model Method ZsRE WikiBio $_{\text{recent}} WikiData$$_{\text{counteract}}$
ES L P HM ES L HM ES L P HM ES L P HM
Llama-2 SERAC 94.9 44.9 40.6 52.2 98.4 70.1 81.9 92.0 33.6 47.0 48.4 78.0 29.3 42.7 42.6
IKE 99.5 46.9 67.6 65.0 99.7 41.6 58.7 85.0 45.8 57.3 58.8 85.7 51.5 60.6 63.0
EREN 36.7 67.1 47.8 47.6 60.8 64.0 62.4 44.4 59.8 38.5 46.0 25.6 67.8 25.4 32.2
WISE 71.1 67.5 51.1 61.9 95.5 99.2 97.3 80.1 48.9 53.2 58.0 72.9 56.6 50.7 58.7
RECIPE 96.4 60.0 52.6 65.1 63.8 61.9 62.8 80.8 61.2 51.4 62.2 55.6 66.8 36.8 49.9
LTE 99.6 49.2 69.2 66.9 100.0 47.2 64.1 100.0 55.0 74.6 72.1 98.4 63.6 73.6 76.0
KEDAS 99.6 91.0 73.6 86.7 100.0 100.0 100.0 100.0 77.6 75.9 83.1 98.4 76.3 77.1 82.8
Qwen2.5 IKE 96.7 42.2 65.1 60.7 99.3 33.4 50.0 93.6 40.7 61.7 58.3 86.2 48.4 59.3 61.1
EREN 34.4 60.7 46.0 44.6 58.1 53.0 55.4 42.1 55.6 37.4 43.8 21.5 62.4 22.0 27.8
WISE 53.0 14.7 25.1 23.7 86.6 100.0 92.8 54.0 96.8 33.0 50.7 45.4 17.9 30.3 27.1
LTE 99.1 48.0 65.4 64.9 98.4 38.7 55.5 99.7 42.5 68.9 62.4 98.0 49.4 69.9 67.0
KEDAS 99.1 89.6 70.2 84.5 98.4 100.0 99.2 99.7 67.1 70.5 76.7 98.0 66.6 73.4 77.2
Table 11: Main results of sequential editing. Metrics include edit success (ES), locality (L), portability (P) and harmonic mean (HM). The highest scores are bolded.
The sequential editing results with Llama-2-7B-Chat Touvron et al. (2023) and Qwen2.5-7B-Instruct Qwen (2025) are presented in Table 11. KEDAS secures the highest harmonic mean scores on all the four datasets, outperforming previous methods significantly.